OpenStudy (anonymous):

Find the product of sq rt of 30 and 6 to the sq rt of 10 and . Express it in standard form sorry

OpenStudy (anonymous):

\[\sqrt[10]{\sqrt{30}\sqrt{6}}\] You mean this?

OpenStudy (anonymous):

\[\sqrt{?}\]